You can afford a $1350 per month mortgage payment. You've found a 30 year loan at 8% interest.

a) How big of a loan can you afford?

$

b) How much total money will you pay the loan company?

$

c) How much of that money is interest?

$
Answer #1
a) P = Regular Payments
PV = Loan Amount
r = rate of interest
n = no of periods
P = r (PV)
1 - (1 + r )-n
1350 = (8%/12)*PV
1 - (1 / (1 + 8%/12)^360))
1350 = 0.00666666666666667 * PV
0.908556628
PV = 1350 * 0.908556627559753 / 0.00666666666666667
PV = 183982.72
b) Total Money Payment = EMI * No of Payment
1350 * 360
486000
c) Interest = Total Payments - Loan
486000 - 183982.72
302017.28